OpenStudy (anonymous):

find the solution set for this equation -b^2-5b=0

OpenStudy (phi):

I would first multiply both sides (and all terms) by -1 that way you get b^2 +5b=0 now factor "b" out of each term. can you do that ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

you should get b(b+5)=0

OpenStudy (anonymous):

yes so

OpenStudy (anonymous):

what do i do next

OpenStudy (phi):

if you multiply two terms and you get zero, that means one or the other must be 0

OpenStudy (anonymous):

so it is (-5,0)

OpenStudy (phi):

you know b*(b+5)=0 first case: b is zero. b=0 second case: b+5=0 to find b, add -5 to both sides to get b+5-5= 0-5 or after simplifying b= -5 so b can be 0 or -5

OpenStudy (anonymous):

thank you it is correct